# Wigner Cat Phases: A finely tunable system for exploring the transition to quantum chaos

### v1

@@ Abstract
The transition to chaos for quantum dynamics is quantified via a finely tunable mixed random matrix ensemble. The mixed Gaussian Orthogonal Ensemble (mGOE) forms a pedagogically accessible family of systems in simulating Many-Body Localization (MBL) transitions. It can be tuned from chaotic to localized and heavy-tailed localized phases in a continuous fashion, providing an opportunity to explore new phases. We numerically study how the spectral properties of mGOE evolve during these transitions. Characterization of transition to quantum chaos is computed and analyzed via empirical spectral density, nearest-neighbor spacing, and adjacent gap ratios with statistical uncertainty quantifications that strengthens the robustness of evidence of transitions. The transition is identified as Wigner Cat Phases, because of the shape of empirical spectral densities, which depens on the tuneable parameter. These simulated phases in mGOE appear to be an ideal tool to study Eigenstate Thermalization Hypothesis (ETH) and its related transitions, representing a family of physical systems under different localisation and disorder strengths.
@@
quantum chaos, spectral fluctuations, random matrix theory, eigenstate thermalization, many-body localisation, quantum thermodynamics
@@ 1 Introduction
Classical to quantum mechanical correspondence in describing how classically chaotic systems could manifest in quantum mechanics has been an interest with great fascination [berry87, ford92, gutzwiller92, heller93]. In this direction, how to define and quantify quantum chaos is extensively studied [berry77, chirikov88, zurek95, chirikov95, stockmann09, gutzwiller13, haake18]. Among these studies, Bohigas-Giannoni-Schmit (BGS) conjecture [bohigas84] took a prominent place due to its clear practical demonstration of a conceptual foundation that invariant random matrix ensembles of Wigner-Dyson [wishart28, wigner51, wigner55, wigner57, wigner58, wigner67, dyson62a, dyson62b, dyson62c, dyson62d, dyson62e, mehta04, tao, potters20] represents quantum systems whereby their classical counterparts are chaotic in Lyapunov sense and ergodic. This notion is also called quantum chaology due to Sir Berry, but we use the term quantum chaos. In a prior work, Sir Berry’s conjecture[berry77, jarzynski97] made a closely related assertion, that excited eigenfunctions are selected randomly from an ensemble of Gaussian functions.
@@
Researchers become aware of the connections of BGS conjecture and how statistical quantum mechanics could satisfy quantum ergodicity by applying random matrix theory perturbations, known as Eigenstate Thermalization Hypothesis (ETH) [jdeutsch91, srednicki94, srednicki94b, srednicki95, srednicki96, srednicki99], reviewed [dalessio16, deutsch18] and revisited with Loschmidt echo or out of time order correlators (OTOC) measures [foini19, papalardi22, foini25]. The similarity of ETH against BGS is also shown [magan24, weiden25]. In these directions, there has been growing interest for quantum chaos in quantum information theory and its applications [cotler17, magan18, ali20, altland24, borner24, anand24, googleq25].
@@
A localisation phenomenon whereby thermalization wouldn’t be achieved fully or partially is an open area of research, i.e., Many-Body Localisation (MBL) [sierant25] and comprehensive references there-in. The primary open questions are: how MBL phases occur? What are their statistical and physical properties from random matrix perspective? Aiming at this, there are works exploring the approach to ETH via building models and methods for studying these transitions: the transition from integrability to nonintegrability in quantum many-body systems [santos04, chavda14, sierant19, pai20, sierant25].
@@
Figure 1: Spectral densities are numerically identified for different tuning parameters, at (1a) $\mu=0.54$, (1b) $\mu=0.70$, (1c) $\mu=0.88$ and (1d) $\mu=0.98$. These are so-called Wigner Cat Phases due to their M-shaped densities deviating from Wigner’s semi-circle law. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that semicircle law is recovered at small mixtures, i.e., higher $\mu$ values.
@@
We presented a new kind of invariant random matrix ensemble, a mixed version of Wigner-Dyson ensembles. Construction of such ensemble is discussed and along with a principled way of building spectral densities in these mixed ensembles. One of this ensemble, mixed Gaussian Orthogonal Ensemble (mGOE) [suzen20, suzen21] is used in simulating transition to ETH and phases of MBL. Simulations give rise to so-called Wigner Cat Phases due to M-shaped empirical spectral densities. Evolution of the nearest neighbor spacing distribution (NNSD) with unfolding with truncation of spectral outliers [ash14, abu18] and the ratio of two consecutive level spacings (adjcency gap) over a continuous transition that is possible with mGOE are numerically identified. These transitions represent different localisation strengths and MBL phases. Adjcency gaps are preferred way of analysis due to their independence to spectral unfolding procedure [oganesyan07, chavda14]. Sampling from mGOE can be made very finely due to continuous mixture parameter, providing unprecedented resolution to study MBL phases.
@@
In Section 2, we shortly provide a glance of the quantum thermalization and localisation problem. The following Section 3 presents our family of quantum systems represented by invariant mixed Wigner-Dyson ensembles. So-called Wigner Cat Phases are demonstrated with numerical experiments with spectral measures of nearest-neighbour and adjacent ratio level spacings in Section 4. We finally conclude our study in Section 5.
@@
Figure 2: Nearest-neighbour spacing for different $\mu$ tuning are shown, (2a) $\mu=0.54$, (2b) $\mu=0.70$, (2c) $\mu=0.86$, and (2d) $\mu=0.98$. Deviation from Wigner-Dyson distribution at smaller $\mu$ values are demonstrated with lower values indicating heavy-tailed distribution. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that Wigner-Dyson distribution is recovered at small mixtures, i.e., higher $\mu$ values.
@@ 2 Quantum thermalization transition to localisation
A definite quantum state $|\Psi(t)\rangle$ can evolve over time given its eigenstates $|\alpha\rangle$ and eigenenergies $E_{\alpha}$ in an exponential fashion with a corresponding observable linear operator $A$,
$|\Psi(t)\rangle=\sum_{\alpha}C_{\alpha}\exp(-iE_{\alpha}t)|\alpha\rangle$
where $\sum_{\alpha}|C_{\alpha}|^{2}=1$, a normalised state. How this system achieves thermal equilibrium, observable reaches to an equilibrium value given temperature $T_{\alpha}$ is shown [jdeutsch91, srednicki94, srednicki94b, srednicki95, srednicki96, srednicki99] by two conditions, so-called Eigenstate Thermalization Hypothesis (ETH):
@@
    Expectation value $\langle\alpha|A|\alpha\rangle$ is a smooth function of eigenenergies $\Phi(E_{\alpha})$ implying individual eigenstates possess thermal properties, $\langle A\rangle_{T_{\alpha}}=\Phi(E_{\alpha})$, independent of initial values.
@@
    Given transition from $\alpha$ to $\beta$, diagonal elements of $\langle\alpha|A|\beta\rangle$ should be much larger than off-diagonal elements. This is important as to recover thermal properties for individual eigenstates.
This is also related to quantum ergodicity because ensemble average $\langle A\rangle$ is equal to time average of $A$ if thermalization is achieved. This outstanding and insightful description of quantum thermalization due to Deutsch-Srednicki provided a basis for showing that, quantum system whereby their classical counterparts are chaotic in Lyapunov sense satisfies these two conditions of ETH.
@@
If we invoke BGS conjecture, as it is shown recently that BGS has better statistical predictive power due its statistical nature [magan24, weiden25]. Hence, Dyson-Wigner matrices in Gaussian Orthogonal Ensemble (GOE) satisfies the termalisation hypothesis, and also descibes quantum chaotic systems. We concentrate on this aspect, that, quantum chaotic systems can thermalize and if we deviate from this behaviour our system will not reach thermal equilibrium. This phenomenon is called Many-Body Localisation (MBL). A more in-depth theoretical review is presented recently by Sierant-Lewenstein et al. [sierant25] and references therein.
The core problem reduced to the statistical identification of deviation from quantum chaos, under which conditions and systems this would occur. A prominent measure in this direction was introduced by Oganesyan-Huse in a seminal work [oganesyan07], adjacent gap ratios on the spectra in investigating the MBL phenomenon. Given $N$ spectral spacings $\delta_{i}$, after sorting eigenenergies, mean adjacent gap ratio $r$ is computed as follows,
@@
They found that r takes values of $0.3860$ and $0.5295$ from full localisation to full thermalization. This transition points corresponds to Poisson and Wigner-Dyson distributed quantum systems. Our tunable ensemble as a core tool explained in the next section, can simulate this transition from ETH and phases of MBL. Fascinating phenomenon that this transition corresponds to transition from quantum integrability to nonintegrability and also transition from quantum non-ergodicity to ergodicity.
There are some basic differences among the concepts of chaotic, ergodic and integrable systems [sierant25], but here we concentrate on the chaotic quantum dynamics and the localisation phenomenon based on spectral analysis due to BGS.
3 Invariant Mixing Ensembles: Tuning for Localisation
The primary invariant ensemble that represents a quantum chaotic system as as proposed by BGS conjecture is Gaussian Orthogonal Ensemble (GOE). Its numerical construction was clearly demonstrated by the seminal paper of Edelman-Rao [edelman05], among other generalised ensembles. Here, we concentrate on only GOE. A given matrix $G_{1}$ having entries drawn from Gaussian distribution with mean $\mu=0$ and standard deviation $\sigma$. We can sample GOE and obtain a member of ensemble $A^{GOE}(N)$ with the following algebraic equality [edelman05],a $NxN$ square matrix,
@@ 3 Invariant Mixing Ensembles: Tuning for Localisation
    Offdiagonals with the following distributions: $offdiag(A^{GOE})\sim\mathcal{N}(0,\sigma/2)$.
If we draw $M$ matrices, we call the resulting set $GOE$ of size $M$ and with square matrix sizes of $N$, denoted by $GOE(M,N)$. By doing this we can estimate bootstrapped confidence intervals on all computed properties [efron94, davison97]. This kind of confidence intervals is not usually practiced in RMT studies but in our context it is important for investigations of the mixed ensembles due to nature of numerical generation of mixed ensembles.
@@
    There will be $M$ different size random matrices drawned from $GOE(n_{i})$.
@@
    The degree of mixture $\mu$ can be intepreted as success probability in a Binomial process, that satisfied $M=\sum_{i}m_{i}$ and limiting case $\mu=1.0$ exactly.
@@
    The set of $n_{i}$ then can be identified by $M$ Binomial trials with the following parametrisation $Binomial(\mu,N)$.
@@
    Then we drawn matrices from conventional GOE, $M$ times $GOE(n_{i})$, whereby size of matrices comes from the Binomial process.
@@
For example, in building $mGOE(M,N,\mu)$, a simple possible case for $M=10$ for $N=1000$ at $\mu=0.8$ a possible set of draws from Binomial process can be the following $Binomial(1000,0.8)$, $[809,813,809,797,808,789,795,810,807,797]$ are the sizes $n_{i}$ of matrices we should generate from $GOE$. Mean of the sizes is $800.1$, in the large $M$ limit this should be exactly $800$.
3.1 Spectral Periodicity: Degenarcy and inducing localisation
@@
In investigating localisation phenomenon, MBL, intuitively the degree of mixture $\mu$ plays a role of tunable parameter that drives the representative quantum system of $mGOE$ to localisation away from thermalization at $\mu=1.0$ due to expressed degeneracy and randomness in generating the mixed ensemble via Binomial draws. Effect of such directed randomness is also observed in deep learning [suzen25g].
The theoretical justification of this behaviour lies in mismatched spectral analysis depending on the matrix sizes. This can be imagined via average behaviour of different spectral distribution. Hence, this generates a defect in spectral density in sampling $mGOE$, driving the system to localisation with decreasing $\mu$. From a physical stand point, equivalently this is a kind of inducing a disorder in generating mixed ensemble, amplified by the spectral periodicity and variation of sizes in mixture. This is not a finite-size effect, rather conjectured to be an inherent property of the mixed ensemble.
@@ 3.2 Self-consistent spectral unfolding
Unfolding the spectra aimed at removing local fluctuations and making them locally flat on average. It is shown in the literature that computation of spectral nearest-neighbor spacings are sensitive to eigenvalue outliers [abu18]. For this reason, we used interquartile range of eigenvalues without any loss of distributional quality, i.e., perfect agreement with the theoretical distributions in the full thermalization GOE case.
@@
Figure 4: Average (Mean) gap ratios over different mixture $\mu$ strengths. These values inform us how localisation changes. Horizontal lines describe reaching to full ETH and full integrability respectively.

### v2

@@ Abstract
A transition to quantum chaos is characterized by a finely tunable mixed random matrix ensemble. The mixed Gaussian Orthogonal Ensemble (mGOE) offers a pedagogically accessible framework for simulating Many-Body Localization (MBL) transitions. It can be tuned from chaotic to localized and heavy-tailed phases in a continuous fashion, providing an opportunity to explore new phases. We numerically study how the spectral properties of mGOE evolve across these transitions. Characterization of the transition to quantum chaos is computed and analyzed via empirical spectral density, nearest-neighbor spacing, and adjacent gap, with statistical uncertainty quantification that strengthens the robustness of the evidence for transitions. The transitions are identified as the Wigner Cat Phases, due to the shape of empirical spectral density, which depend on the tunable parameter. These simulated phases in mGOE provide a valuable framework to study the Eigenstate Thermalization Hypothesis (ETH) and its related transitions, representing a family of systems with tunable localization and disorder.
@@
quantum chaos, spectral fluctuations, random matrix theory, eigenstate thermalization, many-body localization, quantum thermodynamics
@@ 1 Introduction
Classical to quantum mechanical correspondence in describing how classically chaotic systems could manifest in quantum mechanics has been an interest in great fascination [7, 38, 27, 29, 32]. In this direction, how to define and quantify quantum chaos is extensively studied [6, 13, 74, 41, 12, 59, 30, 72, 31, 33]. Among these studies, Bohigas-Giannoni-Schmit (BGS) conjecture [8] took a prominent place due to its clear practical demonstration of a conceptual foundation that invariant random matrix ensembles of Wigner-Dyson represent quantum systems [73, 71, 67, 68, 69, 70, 20, 22, 23, 21, 39, 40, 64, 45] whereby their classical counterparts are chaotic in Lyapunov sense and ergodic. This notion is also called quantum chaology due to Sir Berry, but we use the term quantum chaos. In a prior work, Sir Berry’s conjecture [6, 34] made a closely related assertion, that excited eigenfunctions are selected randomly from an ensemble of Gaussian functions.
@@
Researchers become aware of the connections between BGS conjecture and how statistical quantum mechanics could satisfy quantum ergodicity by applying random matrix theory perturbations, known as Eigenstate Thermalization Hypothesis (ETH) [18, 54, 55, 56, 57, 58], reviewed [15, 19] and revisited with the Loschmidt echo or Out of Time Order Correlators (OTOC) measures [26, 44, 25]. The similarity between ETH and BGS is also shown recently [36, 66]. In these directions, there has been growing interest for quantum chaos in quantum information theory and its applications [14, 37, 3, 4, 9, 5, 28] and especially novel techniques of detecting it [16].
@@
A localization phenomenon whereby thermalization wouldn’t be achieved fully or partially is an open area of research, i.e., Many-Body Localisation (MBL) [52] and comprehensive references therein. The primary open questions are: how MBL phases occur? What are their statistical and physical properties from a random matrix perspective? Aiming at this, there are works exploring the approach to ETH via building models and methods for studying these transitions: the transition from integrability to non-integrability in quantum many-body systems [48, 11, 53, 43, 52].
Recently, multi-parametric ensembles are studied in order to describe non-ergodic regimes [10, 46, 47, 49, 51, 50] with local spectral and eigenfunction statistics. From a different perspective, compare to multi-parametric ensembles, in this work, we present a new kind of invariant random matrix ensemble, a mixed version of Wigner-Dyson ensemble, not to be confused with Brownian ensemble [49]. With this new approach, we can capture transition to ergodic regime with a single continuous parameter on the single ensemble. Construction of such ensemble is discussed and along with a principled way of building spectral densities in these mixed ensembles. One of these ensembles, mixed Gaussian Orthogonal Ensemble (mGOE) [62, 63] can be used in simulating transition to ETH and phases of MBL. Simulations give rise to so-called Wigner Cat Phases due to M-shaped empirical spectral densities, that can be a test bed in detecting quantum chaos as well [16], bringing in global, topological fingerprint.
@@
Figure 1: Spectral densities are numerically identified for different tuning parameters, at (1a) $\mu=0.54$, (1b) $\mu=0.70$, (1c) $\mu=0.88$ and, (1d) $\mu=0.98$. These are so-called Wigner Cat Phases due to their M-shaped densities deviating from Wigner’s semi-circle law. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals appear as error bars. We see that semicircle law is recovered at small mixtures, i.e., higher $\mu$ values.
@@
The evolution of the nearest neighbour spacing distribution (NNSD) with the varying $\mu$ is numerically identified. The spectra computation has been carried out with self-consistent unfolding and truncation of spectral outliers [2, 1]. The ratio of two consecutive level spacings (adjacency gap) over a continuous transition, that is possible with mGOE is also identified. These transitions represent different localisation strengths and MBL phases. As it is indicated in the literature, the adjacency gaps are preferred way of analysis due to their independence to spectral unfolding procedure [42, 11, 35]. Sampling from mGOE can be made very finely due to continuous mixture parameter, providing unprecedented resolution to study MBL phases.
@@
In Section 2, we shortly provide basic definitions of the quantum thermalization and localisation problem. The following Section 3 presents our family of quantum systems, represented by invariant mixed Wigner-Dyson ensembles. So-called Wigner Cat Phases are demonstrated with numerical experiments involving spectral measures of nearest-neighbour and adjacent ratio level spacings, along with topological shape of the empirical spectral density in Section 4. We finally conclude our study in Section 5.
@@
Figure 2: Nearest-neighbour spacings for different $\mu$ values are shown, (2a) $\mu=0.54$, (2b) $\mu=0.70$, (2c) $\mu=0.86$, and (2d) $\mu=0.98$. Deviation from Wigner-Dyson distribution at smaller $\mu$ values is demonstrated with lower values indicating heavy-tailed distribution. There is no observed full integrability even in small $\mu$ values, i.e., Poisson distribution is not observed. Uncertainties are computed over mGOE ensemble via bootstrapped $95\%$ confidence intervals and appear as error bars. We see that Wigner-Dyson distribution is recovered at larger $\mu$ values.
@@ 2 Quantum thermalization transition to localisation
A definite quantum state $|\Psi(t)\rangle$ can evolve over time given its eigenstates $|\alpha\rangle$ and eigenenergies $E_{\alpha}$ in an exponential fashion with a corresponding observable linear operator $A$:
$|\Psi(t)\rangle=\sum_{\alpha}C_{\alpha}\exp(-iE_{\alpha}t)|\alpha\rangle,$
where $\sum_{\alpha}|C_{\alpha}|^{2}=1$, a normalized state. How this system achieves thermal equilibrium, when the observable attains an equilibrium value at temperature $T$, is shown by two conditions of the so-called Eigenstate Thermalization Hypothesis (ETH) [18, 54, 55, 56, 57, 58]:
@@
    The expectation value $\langle\alpha|A|\alpha\rangle$ is a smooth function of the eigenenergies $\Phi(E_{\alpha})$, implying that individual eigenstates possess thermal properties, $\langle A\rangle_{T}=\Phi(E_{\alpha})$, independent of initial values.
@@
    Given a transition from $\alpha$ to $\beta$, the off-diagonal element $\langle\alpha|A|\beta\rangle$ should be much smaller than the diagonal elements, so that thermal properties emerge for individual eigenstates.
This is also related to quantum ergodicity because ensemble average $\langle A\rangle$ is equal to time average of $A$ if thermalization is achieved. This outstanding and insightful description of quantum thermalization due to Deutsch–Srednicki provided a basis for showing that, quantum system whereby their classical counterparts are chaotic in Lyapunov sense satisfies these two conditions of ETH.
@@
If we invoke BGS conjecture, as it is shown recently that BGS has better statistical predictive power due its statistical nature [36, 66]. Hence, Dyson-Wigner matrices in Gaussian Orthogonal Ensemble (GOE) satisfies the thermalization hypothesis, and also describes quantum chaotic systems. We concentrate on this aspect, that, quantum chaotic systems can thermalize and if we deviate from this behaviour our system will not reach thermal equilibrium. This phenomenon is called Many-Body Localisation (MBL). A more in-depth theoretical review is presented recently by Sierant–Lewenstein et al. [52] and references therein.
The core problem reduced to the statistical identification of deviation from quantum chaos, under which conditions and systems this would occur. A prominent measure in this direction was introduced by Oganesyan-Huse in a seminal work [42], adjacent gap ratios on the spectra in investigating the MBL phenomenon. Given $N$ spectral spacings $\delta_{i}$, after sorting eigenenergies, mean adjacent gap ratio $r$ is computed as follows,
@@
They found that r takes values of $0.3860$ and $0.5295$ from full localization to full thermalization. This transition points corresponds to Poisson and Wigner-Dyson distributed quantum systems. Our tunable ensemble as a core tool explained in the next section, can simulate this transition from ETH and phases of MBL. Fascinating phenomenon that this transition corresponds to transition from quantum integrability to non-integrability and also transition from quantum non-ergodicity to ergodicity.
There are some basic differences among the concepts of chaotic, ergodic and integrable systems [52], but here we concentrate on the chaotic quantum dynamics and the localisation phenomenon based on spectral analysis due to BGS.
3 Invariant Mixing Ensembles: Tuning for Localization
The primary invariant ensemble that represents a quantum chaotic system as as proposed by BGS conjecture is Gaussian Orthogonal Ensemble (GOE). Its numerical construction was clearly demonstrated by the seminal paper of Edelman–Rao [24], among other generalized ensembles. Here, we concentrate on only GOE. A given matrix $G_{1}$ having entries drawn from Gaussian distribution with mean 0 and standard deviation $\sigma$. We can sample GOE and obtain a member of ensemble $A^{GOE}(N)$ with the following algebraic equality [24],a $NxN$ square matrix,
@@ 3 Invariant Mixing Ensembles: Tuning for Localization
    Off-diagonals with the following distributions: $offdiag(A^{GOE})\sim\mathcal{N}(0,\sigma/2)$.
If we draw $M$ matrices, we call the resulting set $GOE$ of size $M$ and with square matrix sizes of $N$, denoted by $GOE(M,N)$. By doing this we can estimate bootstrapped confidence intervals on all computed properties [65, 17]. This kind of confidence intervals is not usually practiced in RMT studies but in our context it is important for investigations of the mixed ensembles due to nature of numerical generation of mixed ensembles.
@@
    There will be $M$ different size random matrices drawn from $GOE(n_{i})$.
@@
    The degree of mixture $\mu$ can be interpreted as success probability in a Binomial process, that satisfied $M=\sum_{i}m_{i}$ and limiting case $\mu=1.0$ exactly.
@@
[…]
@@
    Then we draw matrices from conventional GOE, $M$ times $GOE(n_{i})$, whereby size of matrices comes from the Binomial process.
@@
For example, in building $mGOE(M,N,\mu)$, a simple possible case for $M=10$ for $N=1000$ at $\mu=0.8$ a possible set of draws from Binomial process can be the following Binomial(1000,0.8)(1000,0.8), $[809,813,809,797,808,789,795,810,807,797]$ are the sizes $n_{i}$ of matrices we should generate from $GOE$. Mean of the sizes is $800.1$, in the large $M$ limit this should be exactly $800$.
3.1 Spectral Periodicity: Degeneracy and inducing localisation
@@
In investigating localisation phenomenon, MBL, intuitively the degree of mixture $\mu$ plays a role of tunable parameter that drives the representative quantum system of $mGOE$ to localisation away from thermalization at $\mu=1.0$ due to exhibited degeneracy and randomness in generating the mixed ensemble via Binomial draws. Effect of such directed randomness is also observed in deep learning [61].
The theoretical justification of this behaviour lies in mismatched spectral analysis depending on the matrix sizes. This can be imagined via average behaviour of different spectral densities. Hence, this generates a defect in spectral density in sampling $mGOE$, driving the system to localisation with decreasing $\mu$. From a physical standpoint, equivalently this is a kind of induced disorder in generating mixed ensemble, amplified by the spectral periodicity and variation of sizes in mixture. This is not a finite-size effect, rather conjectured to be an inherent property of the mixed ensemble.
@@ 3.2 Self-consistent spectral unfolding
Unfolding the spectra aimed at removing local fluctuations and making them locally flat on average. It is shown in the literature that computation of spectral nearest-neighbor spacings are sensitive to eigenvalue outliers [1]. For this reason, we used interquartile range of eigenvalues without any loss of distributional quality, i.e., perfect agreement with the theoretical distributions in the full thermalization GOE case.
@@
Figure 4: Average (Mean) gap ratios over different mixture $\mu$ strengths. These values inform us how localisation changes. Horizontal lines describe reaching full ETH and full integrability respectively reported average gap ratio statistics in the literature.